# FADA: Fast Diffusion Avatar Synthesis with Mixed-Supervised Multi-CFG Distillation

### v1

@@ 4.3 Ablation Study
Additionally, by comparing the performance of teacher and student models under different data qualities and scales, it can be demonstrated that mixed supervised distillation actually enhances the model’s utilization of moderate-quality data. As depicted in Tables 2 and 3, when the teacher model is trained with a large-scale moderate-quality dataset, a decline in model performance is observed, whereas the student model exhibits the opposite trend.
@@ 5 Conclusion
In summary, we proposed FADA, a fast diffusion avatar synthesis framework with mixed-supervised multi-CFG distillation. By designing a mixed-supervised loss, we leverage data of varying quality to enhance the robustness of generated results. Furthermore, a learnable token-based multi-CFG condition design is introduced to maintain the correlation between the audio and the generated video in the distilled models. The quantitative and qualitative experiments across multiple datasets show that our balanced setting achieves fast inference and high-fidelity generation ability. Ablation studies prove our proposed methods are effective in FADA.

### v2

@@ 4.3 Ablation Study
Additionally, by comparing the performance of teacher and student models under different data qualities and scales, it can be demonstrated that mixed supervised distillation actually enhances the model’s utilization of moderate-quality data. As depicted in Tables 2 and 3, when the teacher model is trained with a large-scale moderate-quality dataset, a decline in model performance is observed, whereas the student model exhibits the opposite trend. Briefly, only mixed-supervised distillation can take advantage of the moderate-quality data.
@@ 5 Conclusion
In summary, we proposed FADA, a fast diffusion avatar synthesis framework with mixed-supervised multi-CFG distillation. By designing a mixed-supervised loss, we leverage data of varying quality to enhance the robustness of generated results. Furthermore, a learnable token-based multi-CFG condition design is introduced to maintain the correlation between the audio and the generated video in the distilled models. The quantitative and qualitative experiments across multiple datasets show that our balanced setting achieves fast inference and high-fidelity generation ability. Ablation studies prove our proposed methods are effective in FADA. The limitations, future works and ethical concerns of this paper will be discussed in Appendix B, H and I respectively.